# NAAQA: A Neural Architecture for Acoustic Question Answering

### v1

@@ 5.4 Feature Extraction
The Interleaved feature extractor (Figure 3(b)) processes 2D spectrogram with a serie of 1D convolutions. The order of the 1D convolution in each block has a significant impact on performance. When the first 1D convolution in each block is done along the frequency axis (Interleaved Freq.), the network reaches an overall accuracy of 59.6%. It performs especially poorly with questions related to notes (34.2%) and instruments (48.0%). The performance on position questions is also the lowest among all the extractors. A possible explanation relate to the nature of the sounds in the CLEAR dataset. Since these sounds are mostly sustained musical notes, their temporal evolution is rather stable. Therefore, the time dimension does not contain much information related to the identification of individual sounds. The time axis however contains important information relative to the scene as a whole which is exploited by higher level layers (Resblocks) to model the connections between different sounds. Because its stride is greater than 1x1, each 1D convolution downsamples the axis on which it is applied. When the first convolution is a frequency convolution, the frequency axis of the resulting features is downsampled which reduces the useful information that can be captured by the time convolution that follows.
When the order of the convolution is reversed (Interleaved Time), useful information is better captured and the network reaches the best overall accuracy with 90.9%. The most notable improvement is observed with questions related to relative position with an increase of 11.1 percentage points compared to the Parallel extractor.
All configurations evaluated in this section struggle with the concept of counting especially when asked to count the number of different instruments playing in a specific part of the scene. They also have difficulties with questions related to the relative position of the instruments. We further discuss these difficulties in Section 6.1. The Interleaved feature extractor is the one that performs the best and constitute the basis of NAAQA in all subsequent experiments.
5.5 Coordinate Maps
@@
–* Both* Both* Both* 94.1 ±0.62 91.2 ±0.33 90.9 ±0.21
Both 93.3 ±0.18 91.0 ±0.22 90.9 ±0.18
Time 93.7 ±0.49 90.5 ±0.70 90.5 ±0.54
Time 93.6 ±0.81 90.7 ±0.92 90.4 ±0.74
Both 93.5 ±0.59 90.2 ±0.77 90.0 ±0.85
Time 93.6 ±1.08 90.0 ±0.95 89.8 ±1.12
Freq 84.6 ±0.96 76.1 ±3.97 75.7 ±3.74
Both 84.5 ±5.66 76.2 ±3.56 75.2 ±2.29
Freq 84.2 ±2.42 74.7 ±1.33 73.9 ±1.54
81.2 ±4.13 73.4 ±1.14 73.5 ±1.07
Freq 80.5 ±4.82 73.1 ±5.39 70.5 ±5.65
Freq 82.2 ±2.09 72.2 ±1.86 69.7 ±2.27
Time 70.9 ±17.56 63.7 ±13.95 61.8 ±12.63
Both 46.5 ±0.37 45.8 ±0.09 45.7 ±0.10
@@
TABLE IV: Impact of the placement of Time and Frequency coordinate maps. All possible positions are illustrated by the pink border boxes in Figure 2. The value Both indicate that both Time and Frequency coordinate maps were inserted at the given position. The interleaved feature extractor is used with hyper-parameters from the initial configuration (defined in section 5.2). The row marked with * correspond to the placement in the initial configuration.
@@
None 73.5 ±1.07 90.9 ±0.50 84.3 ±0.73 95.8 ±0.32 95.7 ±0.54 84.0 ±1.06 48.1 ±2.40 78.3 ±3.11 48.9 ±4.23 51.7 ±1.65 61.3 ±0.63 29.6 ±1.07
Freq only 70.5 ±5.65 88.2 ±3.36 80.1 ±5.40 94.9 ±1.29 93.7 ±1.69 81.0 ±3.24 43.6 ±11.25 74.4 ±8.31 49.9 ±5.85 50.2 ±6.52 60.4 ±2.40 31.1 ±0.62
Time only 90.4 ±0.74 96.9 ±0.42 93.4 ±0.82 98.2 ±0.12 98.1 ±0.49 88.6 ±0.60 94.9 ±1.00 96.7 ±0.90 69.2 ±4.84 64.2 ±1.38 63.6 ±1.02 39.7 ±1.89
Time & Freq 90.9 ±0.18 97.0 ±0.26 94.0 ±0.14 98.5 ±0.31 97.9 ±0.15 88.8 ±0.32 96.3 ±0.29 97.5 ±0.06 67.1 ±5.47 65.6 ±1.14 62.7 ±1.09 38.9 ±2.50
@@
TABLE V: Comparison of Time and Frequency coordinate maps. Coordinate maps are inserted in the Resblocks. Test accuracy is reported by question type (Examples for each question type are available in Table I).
We analyzed the impact of the placement of Time and Frequency coordinate maps at different depths in the network. The pink border boxes in Figure 2 show all the locations where coordinate maps can be inserted prior to convolution operations. In addition to these locations, the insertion of coordinate maps through the Interleaved feature extractor was also explored. All possible locations were evaluated via grid-search. For each location, we inserted either a Time coordinate map, a Frequency coordinate map or both. Results are detailed in Table IV. Coordinate maps have the biggest impact on performance when inserted in the first convolution after the feature extractor and in the Resblocks. This could be because the fusion of the textual and acoustic features, and therefore most of the reasoning, is performed in the Resblocks. The network might be using the additional localization information to inform the modulation of the convolutional feature maps based on the context of the question.
@@
The impact of the kind of coordinate maps on each question types is further analyzed in Table V. To facilitate comparison, all coordinate maps are inserted at the same location (in the first convolution after the feature extractor). When the network does not include any coordinate maps it reaches only 73.5%. Inserting a Frequency coordinate map does not increase performance. As hypothesized, Time coordinate maps significantly increase the overall accuracy with 90.4%. The accuracy increases for all question types but is most notable with questions related to Absolute position where the accuracy goes from 48.1% to 94.9%. This is a similar result to [undefaam] where coordinate maps were proved to help convolutional networks locate objects in visual scenes. In the context of a spectro-temporal representation, this ability translates to enhanced temporal localization which naturally helps to answer questions related to position. Questions related to Notes also receive a significant boost from Time coordinate maps going from 84.3% to 93.4%. The fact that Frequency coordinate maps have a negligible impact on questions related to Notes is rather surprising since notes are mostly defined by the frequency content at a given point in time. Overall accuracy is slightly increased when combining both Time and Frequency coordinate maps with 90.9%.
5.6 Complexity reduction
One aim of this study is to determine what is the smallest configuration (in terms of number of parameters) that can achieve reasonable performance on this acoustic task. Most parameters in NAAQA are used for the feature modulation in the FiLM layers. The number of feature modulation parameters is proportional to the number of GRU text-processing units $G$. The number of Resblock $J$ dictates the number of FiLM layers in the model and therefore the number of modulation coefficients to compute. The number of convolutional filters $M$ in each block also impacts the modulation calculation. The classifier module, which comprises a $1\times 1$ convolution with $C$ filters and a fully connected layer with $H$ hidden units, accounts for almost 10% of the model parameters. Detailed results of the optimization process and values for all explored hyper-parameters ($G,J,M,C,H$) are reported in appendix A.
The NAAQA’s original configuration described in Section 5.2 comprises 5.61M parameters and achieves 90.9%. The most notable complexity reduction comes from the reduction of the number of GRU units $G$. With $G=512$, the number of parameters is reduced to 1.94M and the accuracy increases to 91.1% showing that similar performance can be achieved with almost 3 times fewer parameters. We found that a model comprising $\sim$19 times fewer parameters than the original configuration (0.30M parameters) can achieves better results than the baseline with 84.7% vs 77.9%. The final NAAQA configuration comprises $\sim$7 times fewer parameters (0.81M) than the original configuration and achieves 91.5%. It is composed of an Interleaved feature extractor with $K=3$ blocks and $P=32$, $G=512$ GRU units, $J=4$ Resblocks with $M=64$ filters, a classifier module with $C=256$ and $H=1024$.
5.7 Impact of dataset composition
We investigated the impact of the dataset composition and, specifically, the importance of the number of questions per scene. In order to eliminate the effect of overall size of the data, multiple versions of the dataset were generated by varying the number of scenes while keeping the total number of records fixed. We experiment with 10k, 20k, 50k, 100k, 200k and 400k scenes ($S$). The number of questions per scene for each configuration is defined by $\frac{\text{D}}{\text{S}}$ where $D$ is the total number of records. We also investigate the impact of the dataset total size by repeating this experiment with D=100k,200k,400kD={100k,200k,400k} records.
The network is trained on each of these sets using 70% of the data. To facilitate comparison, the model is evaluated on the same test set as in previous experiments. Results are in Figure 4.
The total number of training records is the most determining factor for the test accuracy. All experiments trained on 400k records perform better than the ones on 200k records and the same for the one trained on 200k records versus 100k records. This is unsurprising since more training examples often lead to better performance. We also expected that having more questions for each scene in the training set would increase the network performance. The results however show that the number of questions per scene did not substantially affect accuracy, especially with larger sets.
Figure 4: Test accuracy for NAAQA trained on different dataset sizes. The total number of records is either 100k, 200k or 400k. The $x$-axis is the number of scenes and the number of questions per scene used to build the dataset. The $y$-axis is the accuracy on the test set. The horizontal line is the mean accuracy. The dataset configuration highlighted in green is the configuration used for all previous experiments.
6 Discussion
Figure 5: Test accuracy of NAAQA final configuration by question type and the number of relation in the question. The presence of before or after in a question constitute a temporal relation. The accuracy is N/A for relative position and count compare since these types of question do no include relations. The hyper-parameters used are described at the end of Section 5.6
6.1 NAAQA detailed analysis
NAAQA performs well on this AQA task with 91.6% accuracy. It does however struggle with certain types of question as shown in Figure 5. When asked to count the number of sounds with specific attributes, NAAQA reaches only 63.3% accuracy. It attains similar accuracy when asked to compare the number of instances of acoustic objects (more, fewer or equal number) with specific attributes (63.5%). This shows that the network learned to compare quantity but cannot achieves better accuracy because of its limited ability to count. The network can successfully recognize individual instruments in the scene (96.1%) but cannot count the number of different instruments playing in a given part of the scene (39.1%). This type of question is rather complex to answer. Consider for example the question : ”How many different instruments are playing after the third cello playing a C# note?”. The model has to first find the cello playing the C# note, then identify all acoustic objects that are playing after this sound, determine which instruments are of the same family and finally count the number of different family. This requires the network to focus on many different acoustic objects. In comparison, the network only needs to focus on 1 objects when answering a question like ”What is the note played by the dark loud violin?”. The same applies for instrument, brightness, loudness and exist questions. The difficulty of focusing on multiple acoustic object in the scene might be an explanation as to why NAAQA struggles with counting. It could also explain why the model struggles with questions related to the relative position of an instrument (66.7%). To answer the question ”Among the flute sounds, which one plays an F note?”, the model must find all the flutes playing in the scene, determine which one play an F note, count the number of flute playing before and translate the count to a position [^5] This is only one possible strategy to answer the question. There may be other ways.. This also requires the network to focus on multiple acoustic objects.
Certain questions include temporal relations between sounds (before and after) as exemplified in Table I. The number of relations in a question correlates with the number of sounds that need to be focused on to answer a question. Figure 5 shows the accuracy based on the number of relations for each question types. Questions that require the network to focus on a single acoustic object (brightness, loudness, instrument, note, global position and absolute position) benefit from the presence of a relation in the question. This might be explained by the fact that the question contains more information about the scene which helps to focus on the right acoustic object. However, the presence of relation in questions that already require the network to focus on multiple objects (exist, count and count comparison) is detrimental. This again supports the idea that having to focus on too many objects in the scene hinders the network’s performance.
6.2 Importance of text versus audio modality
In Section 5.7, we observed that a higher number of questions per scene during training did not improve the test accuracy. This contradicted our hypothesis that increasing the number of questions per scene would increase the performance of the model. It also suggests that the network uses both the scene and question to solve the task without favoring one or the other. To validate the fact that NAAQA makes good use of both input modalities (spectrogram and text), the network was also trained using only one modality. To do so, either all values of the input spectrograms were set to 1 or all questions were set to the unknown token during training. No more than 48% of accuracy on the test set was achieved for either of these scenarios which indicates that the network uses information from both modalities to achieve high performance and that the dataset cannot be solved using only one modality with this architecture. This is however much higher than the probability of answering the correct answer by chance ($\frac{1}{57}=1.75\%$), or by majority vote (7.5%) introduced in Section 3. Although we have carefully designed the data set in order to avoid bias, some form of bias is intrinsic in QA problems. For example, an acoustic sequence containing only $n$ elementary sounds automatically disables all counting answers above $n$ regardless which question is asked. Similarly, asking a yes/no question limits the number of possible answers to two, regardless of the acoustic scene. The denominator used in the chance level (57) will, therefore be lower on average. It remains to investigate if our results are consistent with this argument, or if they suggest a residual bias in the dataset. We leave this investigation for future work.
6.3 Variability of the input size
@@
Batch normalization has been proven to help regularize the training of neural networks [undefaan]. However, using batch normalization requires all inputs in a batch to have the same dimensions in order to compute the element-wise means and standard deviations. Acoustic signals in the CLEAR dataset have variable duration. To work around this constraint, we padded the spectrograms with zeros so that they all have the same dimensions. Another solution is to resize all spectrograms to a fixed size using bilinear interpolation as it is commonly done in the visual domain. From a purely acoustic point of view, we hypothesized that padding spectrograms would preserve the relative time axis between all scenes and would yield better results than resizing. We revisited this hypothesis by training our best architecture on resized spectrograms. Surprisingly, the network performed better by about 1 percentage points when trained on the resized spectrograms. We do not have a definitive answer as to why we observe this behavior. It might be due to the fact that padded spectrograms contain less information since a portion of them is filled with zeros. The network could also be using the different time resolutions in the resized spectrograms as a cue to cheat the reasoning process. However, this could also happen when the scenes are padded since the padded section is unique to each scene. We analyzed the network activations and did not find that the padded section was overly activated. We leave it up to future work to further analyze this.
7 Conclusions
In this study, we propose a version of the CLEAR dataset that emphasizes the specific challenges in acoustic question answering by introducing scenes of variable duration and a greater variety of elementary sounds. We showed that Visual FiLM [undefab], which is a neural network designed for visual question answering, achieves relatively good accuracy on this acoustic task although it struggles with certain types of questions. We propose the NAAQA architecture where the feature extraction is optimized to the acoustic task. This Interleaved feature extractor makes use of 1D convolutions to independently process each axis of 2D time-frequency representations. The initial NAAQA architecture reaches 90.9% of accuracy on the CLEAR task with 5.61M parameters. The final NAAQA configuration reaches 91.5% of accuracy with $\sim$7 times fewer parameters (0.81M). The effectiveness of coordinate maps in this acoustic context was also studied. We showed that time coordinate maps enhance the network temporal localization capabilities which translates in greater performance on this task. We hope that these results and the software we provide will stimulate research in acoustic reasoning.
Acknowledgment
We would like to acknowledge the NVIDIA Corporation for donating a number of GPUs. Part of this research was financed by the CHIST-ERA IGLU project, the CRSNG, the Michael-Smith scholarships and the University of Sherbrooke.
Appendix A Network Optimization

### v2

@@ 6.1.1 Feature Extraction
We then tested two versions of the Interleaved feature extractor (Figure 3(b)). Results show that the order of the 1D convolutions in each block has a significant impact on performance. When the first 1D convolution in each block is done along the frequency axis (NAAQA Interleaved Freq), the network reaches an overall accuracy of 67.2%. It performs especially poorly with questions related to count (42.5%), count instruments (46.5%) and notes (48.0%). The performance on position questions is also the lowest among all the extractors. When the order of the convolution is reversed (NAAQA Interleaved Time), useful information is better captured and the network reaches 78.0% of overall accuracy. A possible explanation relates to the nature of the sounds in the CLEAR2 dataset which mainly consists of sustained musical notes. The time dimension at short scales does not contain much information that helps identifying the individual sounds. At larger scales, however, the time axis contains information relative to the scene as a whole which is exploited by higher level layers (Resblocks) to model the connections between different sounds. Because its stride is greater than 1x1, each 1D convolution downsamples the axis on which it is applied. When the first convolution is a frequency convolution, the frequency axis of the resulting features is downsampled which reduces the useful information that can be captured by the time convolution that follows.
The Parallel feature extractor (NAAQA Parallel, Figure 3(a)) reaches an overall accuracy of 78.5%. It performs well on all question’s types except relative position, count and count instrument. Refer to Section 6.2 for further analysis. These results show that building complex time and frequency feature separately and fusing them at a later stage is a good strategy to learn acoustic features for this task. This claim is further strengthen by the analysis of section 6.3.2.
Out of all extractors, NAAQA Parallel is the one that performs the best and constitutes the basis of NAAQA in all subsequent experiments.
6.1.2 Coordinate Maps
@@
Time 95.0 ±0.79 90.2 ±0.62 79.0 ±0.44
Time 95.1 ±0.90 90.0 ±0.28 79.0 ±0.43
Both 94.9 ±1.08 90.1 ±0.83 78.8 ±0.52
Both 95.1 ±1.23 90.0 ±0.46 78.7 ±0.70
Time 94.7 ±1.06 88.3 ±1.24 76.7 ±0.83
Both 94.1 ±1.47 88.4 ±1.06 76.6 ±0.36
Freq 84.0 ±2.17 73.5 ±1.40 64.8 ±1.51
85.2 ±0.71 72.9 ±1.36 63.8 ±0.70
Both 83.5 ±3.46 72.3 ±2.71 63.5 ±1.93
Freq 84.5 ±0.67 71.7 ±2.85 62.6 ±2.36
Freq 83.4 ±1.53 70.7 ±1.74 61.7 ±1.16
Time 81.9 ±0.88 70.1 ±1.99 61.7 ±0.76
Freq 79.7 ±3.70 67.1 ±3.36 59.3 ±2.20
@@
TABLE V: Impact of the placement of Time and Frequency coordinate maps. All possible positions are illustrated by the pink border boxes in Figure 2. The value Both indicate that both Time and Frequency coordinate maps were inserted at the given position. The NAAQA Parallel is used with hyper-parameters from the initial configuration (defined in section 5.2. The rows are ordered by test accuracy.).
Coordinate maps can be inserted before any convolution operation (Figure 2). We therefore analyzed the impact of the placement of Time and Frequency coordinate maps at different depths in the network. All possible locations were evaluated via grid-search. For each location, we inserted either a Time coordinate map, a Frequency coordinate map or both. Results are detailed in Table V. Time coordinate maps have the biggest impact on performance, especially when inserted in the first convolution after the feature extractor or in the Resblocks. This could be because the fusion of the textual and acoustic features, and therefore most of the reasoning, is performed in the Resblocks. The network might be using the additional localization information to inform the modulation of the convolutional feature maps based on the context of the question. Surprisingly, the Frequency coordinate maps have a minimal impact on performance. We further compare the impact of Time versus Frequency coordinate maps in Supplementary B.
6.1.3 Complexity Optimization
Figure 4: Test accuracy by question type and the number of relation in the question for Optimized NAAQA Parallel. The overall accuracy for this configuration is 79.1%. The presence of before or after in a question constitutes a temporal relation. The accuracy is N/A for relative position and count compare since these types of question do no include relations. The hyper-parameters used are described at the end of Section 6.1.3
As described in Section 4.2.3, we optimized the most important hyper-parameters ($G,J,M$) in the NAAQA model in order to reduce its complexity. The baseline Visual FiLM Resnet101 configuration comprises 6.71 M parameters and achieves only 62.3%. NAAQA Parallel comprises 5.61 M parameters and performs significantly better with 78.5%. With this model as starting point, we performed an ablation study to find which hyper-parameters can be reduced without impacting accuracy. The Optimized NAAQA Parallel configuration is the best trade-off between model complexity and performance. It comprises 1.68 M parameters and achieves the best overall accuracy with 79.5%. The most notable complexity reduction comes from the reduction of the number of GRU units $G$. Reducing $G$ from $4096$ to $512$ increased accuracy while reducing the number of parameters by a factor of 3 (6.61 M vs 1.68 M). The Optimized NAAQA Parallel is composed of a Parallel extractor with $K=3$ blocks and $P=64$, $G=512$ GRU units, $J=4$ Resblocks with $M=128$ filters, a classifier module with $C=512$ filters and $H=1024$ units. Results for this configuration can be found in Table III and Figure 4. Further results related to the ablation study can be found in Supplementary H.
6.1.4 Adding a MALiMo controller
The bottom rows of Table III show results where the configurations described in previous sections are augmented with a MALiMo controller. Although the model complexity is significantly increased ($\sim$1M parameters), this addition does not bring any improvement in the model performance on CLEAR2. Almost all the tested configurations with a MALiMo controller perform slightly worse than the same configuration without the module, as can be seen in Table III. This may be again explained by the characteristics of the sounds in CLEAR2. A more in-depth discussion is given when we evaluate the models on DAQA.
6.2 Summary of Results on CLEAR2
NAAQA performs well on the CLEAR2 AQA task with 79.5% overall accuracy. It does however struggle with certain types of question as shown in Table III and Figure 4. When asked to count the number of sounds with specific attributes, NAAQA reaches only 53.8% accuracy. This limitation is more severe if the question is to count the different instruments playing in a given part of the scene (50.4%). It attains slightly higher accuracy when asked to compare the number of instances of acoustic objects (more, fewer or equal number) with specific attributes (60.6%). In contrast, the network can successfully recognize individual instruments in the scene (81.7%). This suggests, that the problem lies in the logical complexity of the question rather than in the pattern matching from the acoustic scene. As an example, the question (count instrument): ”How many different instruments are playing after the third cello playing a C# note?” requires to first identify the cello playing the C# note, then identify all acoustic objects that are playing after this sound, determine which instruments are of the same family and finally count the number of different families. The model struggles when it must focus on a large number of acoustic objects which explain the low accuracy for this type of question.
A similar argument could explain why models also have difficulties with questions related to the relative position of the instruments (58.0%). For example, to answer the question ”Among the flute sounds, which one plays an F note?”, the model must find all the flutes playing in the scene, determine which one play an F note, count the number of flute playing before and translate the count to a position [^4] This is only one possible strategy to answer the question. There may be other ways.. This also requires the network to focus on multiple objects.
Certain questions include temporal relations between sounds (before and after) as exemplified in Table I. Questions that include relations require focusing on several sounds to be answered. Figure 4 shows the accuracy for each question type depending on the presence of temporal relations. Questions that require the network to focus on a single acoustic object (brightness, loudness, instrument, note, global position and absolute position) benefit from the presence of a relation in the question. This might be explained by the fact that the question contains more information about the scene which helps to focus on the right acoustic object. However, the presence of relations in questions that already require the network to focus on multiple objects (exist, count and count comparison) is detrimental. This again supports the idea that having to focus on too many objects in the scene hinders the network performance.
6.3 Evaluation of DAQA
In order to compare our results to those of [undefaa] [undefaa], we evaluated our models on a version of the DAQA data set. As mentioned in Section 3.2, we were not able to reproduce the original DAQA dataset which mean that the results presented in this section are not fully comparable with [undefaa]. We report results for different configurations of NAAQA on our modified DAQA in Table IV.
6.3.1 NAAQA on DAQA
The models explored in this section are considerably smaller in terms of number of parameters than the ones explored in [undefaa]. The smallest model they evaluated had 5.49M parameters, the biggest model had 21.33M parameters and the best performing model had 13.20M parameters. The Optimized NAAQA 2D Conv model only has 1.68 M parameters and reaches an accuracy of 58.3% on DAQA. The Optimized NAAQA Parallel has the same number of parameters and performs slightly better with and accuracy of 60.4%. When we analyzed both of these models on CLEAR2 dataset in Section 6.1.1, we found a much smaller difference between the performance of the NAAQA 2D Conv and the NAAQA Parallel. This difference in the observations suggests that the parallel extractor is more useful in the context of complex acoustic sounds (DAQA) than with sustained musical note (CLEAR2).
Even though these results are not 100% comparable with [undefaa] because of the difference in the dataset composition, we want to emphasize that Optimized NAAQA Parallel reaches a somewhat similar accuracy than the smallest FiLM in [undefaa] (60.4% vs 64.3%) with significantly smaller number of parameters (1.68 M vs 5.49 M).
6.3.2 NAAQA with a MALiMo module on DAQA
In Section 6.1.4, we found that adding a MALiMo controller to our NAAQA models did not improve the accuracy on CLEAR2. Results in Table IV show that the MALiMo controller have a significant impact on performance when the model is evaluated on DAQA dataset. We see an increase of almost 4% when using Optimized NAAQA Parallel + MALiMo ctrl compared to Optimized NAAQA Parallel alone. These results are consistent with [undefaa] findings and with the hypothesis that MALiMo increases performance when working with complex sounds.
The Optimized NAAQA Parallel + MALiMo ctrl configuration performs about the same as the smallest MALiMo model evaluated in [undefaa] (64.3% vs 65.1%) with significantly less parameters (2.78 M vs 8.91 M).
7 Conclusions
Acoustic Question Answering (AQA) is a newly emerging task in the area of machine learning research. As performance is strongly dependent on the acoustical environments and types of questions, it is important to understand the relationship between the application and the chosen neural architecture. We propose a benchmark for AQA based on musical sounds (CLEAR2) and a neural architecture that is tailored to interpreting acoustic scenes (NAAQA). NAAQA introduces a number of modifications to a FiLM based architecture in order to optimize acoustic scenes analysis. These includes several strategies for neural feature extraction, an ablation study of the hyper-parameters and the optimization of coordinate maps for this task. We confirm that FiLM layers are very effective to modulate activation maps in the AQA application. We are able to optimize our NAAQA neural network so to obtain competitive results with a fraction of the model complexity. These results are confirmed on a different AQA task (DAQA) comprising more complex sounds with the addition of a MALiMo controller in the model. We release all code openly in the hope that these resources may foster increased research activity in solving the AQA task.
Acknowledgment
We would like to acknowledge the NVIDIA Corporation for donating a number of GPUs. Part of this research was funded or supported by the CHIST-ERA IGLU project, the CRSNG, the Michael-Smith scholarships the University of Sherbrooke and NTNU.
Supplementary Material
Supplementary A Detailed training settings
All models were trained for a maximum of 40 epochs or until the validation loss stops decreasing for 6 consecutive epochs. The Adam optimizer [undefaav] with a learning rate of $3\text{\times}{10}^{-4}$ was used to optimize the cross entropy loss. To reduce the likelihood of the model getting trapped in a local minima, the learning rate is automatically reduced by a 0.1 factor if the validation loss stops decreasing for 3 consecutive epochs. L2 regularization with a weight decay of $5\text{\times}{10}^{-6}$, dropout [undefaaw] with a drop probability of 0.25 and batch normalization [undefaap] are used to regularize the learning process. A batch size of 128 is used for all experiments. The following 5 random seeds were used to run all experiments : 876944 189369 682421 175326 427438
Supplementary B Time and Frequency coordinate maps impact
@@ Acknowledgment
Time only 79.0 ±0.43 80.8 ±1.22 73.0 ±0.84 90.8 ±0.52 86.8 ±0.49 80.2 ±0.75 80.2 ±0.89 89.0 ±0.60 53.1 ±0.90 52.1 ±0.70 61.1 ±0.56 49.3 ±0.47
Time & Freq 78.7 ±0.70 80.3 ±0.92 72.5 ±0.82 90.8 ±0.83 87.0 ±1.02 80.6 ±0.47 79.9 ±1.25 88.9 ±0.58 54.8 ±1.21 52.9 ±0.42 59.1 ±2.46 46.0 ±3.14
None 63.8 ±0.70 73.3 ±0.72 61.8 ±0.62 87.6 ±0.66 84.6 ±0.96 77.1 ±0.56 41.0 ±2.19 69.8 ±1.33 49.0 ±2.02 45.9 ±0.80 59.3 ±0.41 44.6 ±1.71
Freq only 61.7 ±1.16 72.7 ±0.88 60.4 ±0.89 87.4 ±0.62 83.4 ±0.41 75.6 ±0.71 36.8 ±3.80 67.2 ±1.58 47.6 ±1.86 44.2 ±1.58 57.0 ±2.80 43.5 ±4.09
@@
TABLE VI: Comparison of Time and Frequency coordinate maps. Coordinate maps are inserted in the Resblocks. Test accuracy is reported by question type (Examples for each question type are available in Table I).
@@
The impact of the kind of coordinate maps on each question types is presented in Table VI. To facilitate comparison, all coordinate maps are inserted at the same location (in the first convolution after the feature extractor). When the network does not include any coordinate maps it reaches only 63.8%. Inserting a Frequency coordinate map decreases accuracy to 61.7%. As hypothesized, Time coordinate maps significantly increase the overall accuracy with 79%. The accuracy increase for all question types is most notable with questions related to Absolute position where the accuracy goes from 41.0% to 80.2%. This is a similar result to [undefaaj] where coordinate maps were proved to help convolutional networks locate objects in visual scenes. In the context of a spectro-temporal representation, this ability translates to enhanced temporal localization which naturally helps to answer questions related to position. Questions related to Notes also receive a significant boost from Time coordinate maps going from 61.8% to 73%. The fact that Frequency coordinate maps decrease performance on questions related to Notes is rather surprising since notes are mostly defined by the frequency content at a given point in time. Furthermore, combining both Time and Frequency coordinate maps slightly reduces the overall performance when compared to using only Time coordinate maps which is in line with previous results.
Supplementary C Results showing Effects of stride size on both CLEAR2 and DAQA’
Model Nb Parameters Train acc Val acc Test acc Trainig time
On CLEAR2 dataset
Optimized NAAQA Parallel- Long stride 1.68 M 95.2 ±0.43 90.5 ±0.28 79.1 ±0.32 0 days 02:39:51
Optimized NAAQA Parallel- Short stride 1.68 M 93.1 ±0.48 89.5 ±0.38 77.9 ±0.68 0 days 07:00:01
On DAQA dataset
Optimized NAAQA Parallel- Short Stride 1.68 M 68.0 ±0.22 64.6 ±0.34 65.3 ±0.37 3 days 17:27:45
Optimized NAAQA Parallel- Long Stride 1.68 M 67.3 ±1.38 60.2 ±0.24 60.3 ±0.20 0 days 07:14:17
TABLE VII: Short stride and Long Stride comparison The optimized NAAQA on CLEAR2 and DAQA’. The rows are ordered by test accuracy.
Supplementary D Discussion on results with Resnet
Multiple feature extractors were evaluated in Section 6.1.1. One of them is the Resnet-101 pre-trained on ImageNet (Visual dataset). This extractor achieves the lowest of all tested extractors with 62.3%. However, it performs quite well with questions related to brightness and loudness with 83.9% and 81.9% respectively. One possible explanation is that both brightness and loudness are directly related to the amplitude of the “pixels” along the frequency axis which may be efficiently encoded by the Resnet101 model. The baseline does not perform as well with questions related to absolute position which is somewhat surprising since these questions correspond to a similar pattern matching problem on the time axis instead of the frequency axis. The baseline model also has difficulties with questions related to notes (37.8%). This might be explained by the fact that a note can be identified by its fundamental frequency and harmonics which can be far apart on the frequency axis. Visual models are trained to recognize localized features and therefore struggle to recognize notes. This is a feature that is typical of acoustic signals.
It does however perform better with global position questions. This is logical because this type of question refers to an approximate position (beginning, middle, end) which constrain the number of possible answers to 3 instead of 15.
Supplementary E Variability of the input size
@@
Batch normalization has been proven to help regularize the training of neural networks [undefaap]. However, using batch normalization requires all inputs in a batch to have the same dimensions in order to compute the element-wise means and standard deviations. Acoustic signals in the CLEAR dataset have variable duration. To work around this constraint, we padded the spectrograms with zeros so that they all have the same dimensions. Another solution is to resize all spectrograms to a fixed size using bilinear interpolation as it is commonly done in the visual domain. From a purely acoustic point of view, we hypothesized that padding spectrograms would preserve the relative time axis between all scenes and would yield better results than resizing. We revisited this hypothesis by training our best architecture on resized spectrograms. Surprisingly, the network performed better by about 1 percentage points when trained on the resized spectrograms. We do not have a definitive answer as to why we observe this behavior. It might be due to the fact that padded spectrograms contain less information since a portion of them is filled with zeros. The network could also be using the different time resolutions in the resized spectrograms as a cue to cheat the reasoning process. However, this could also happen when the scenes are padded since the padded section is unique to each scene. We analyzed the network activations and did not find that the padded section was overly activated. We leave it up to future work to further analyze this.
Supplementary F Importance of text versus audio modality
To validate the fact that NAAQA makes good use of both input modalities (spectrogram and text), the network was trained using only one modality. To do so, either all values of the input spectrograms were set to 1 or all questions were set to the unknown token during training. No more than 48% of accuracy on the test set was achieved for either of these scenarios which indicates that the network uses information from both modalities to achieve high performance and that the dataset cannot be solved using only one modality with this architecture. This is however much higher than the probability of answering the correct answer by chance ($\frac{1}{57}=1.75\%$), or by majority vote (7.5%) introduced in Section 3.1. Although we have carefully designed the data set in order to avoid bias, some form of bias is intrinsic in QA problems. For example, an acoustic sequence containing only $n$ elementary sounds automatically disables all counting answers above $n$ regardless which question is asked. Similarly, asking a yes/no question limits the number of possible answers to two, regardless of the acoustic scene.
Supplementary G DAQA’ missing sounds
Here is a list of all the missing elementary sounds that were originally in DAQA and that are not in the DAQA’ dataset (because of not being anymore accessible on youtube) :
b000_20, b000_3, b000_4, b000_5, b000_6, b001_11, c001_13, c001_14, c001_15, c001_16, c001_17, c001_18, c001_19, c001_20, c003_9, c004_10, d001_1, d001_9, d002_19, f001_19, h001_10, h001_11, h001_12, h001_13, h001_14, h001_15, h001_19, h001_2, h001_3, h001_4, h001_5, h001_6, h001_7, h001_8, h001_9, h002_10, h002_11, h002_12, h002_8, h002_9, t000_14, t000_15
Supplementary H Network Optimization